# Catalog of Mock Stellar Streams in Milky Way-like Galaxies

### v1

@@ 2.3 Stellar Stream Generation
In this section, we describe the complete process for stream generation, given the Galactic potential and our initial GC conditions from the model of Chen & Gnedin (2024a). The distinct inputs to the model are as follows:
@@ 2.3.1 Mass Loss Calculation
The overall mass-loss calculation is in two parts. From cluster birth until $z=0.3$, we use the mass loss given by the GC formation model. Then beginning at $z=0.3$, we numerically solve Equation 2 given the tidal tensor along cluster orbit computed with agama, to calculate the amount of mass lost by the cluster from $z=0.3$ until the present day. While agama uses an adaptive timestep for orbit integration, we interpolate the solution such that the mass history and tidal tensor components are provided at 1000 uniformly spaced intervals over the integration time, giving us both quantities with a time resolution of $3.5$ Myr. Figure 12 in Appendix A shows the effective tidal frequency as a function of time at the provided resolution for an example cluster in one of the TNG halos. There are only four full TNG snapshots with potential information in our chosen time of integration, which miss significant variation of the tidal force. The agama framework is much more accurate for detailed calculations of the mass loss.
@@ 2.3.2 Stream Realization
Once calculated, the mass-loss histories are passed onto the next step, where we initialize a stellar population for each cluster and compute the number and masses of stream stars to be released at each timestep. First, individual cluster stars are sampled according to a Kroupa (2001) IMF with the initial cluster mass given by the GC model. For the lower and upper bounds of the IMF, we use MIST isochrones (Choi et al., 2016; Dotter, 2016) to calculate the theoretical lowest and highest mass stars that would still be luminous today given the cluster’s age and metallicity. These values vary from cluster to cluster but are generally around $0.1\,{\mathrm{M}_{\odot}}$ for the lower bound and between $0.8-0.9\,{\mathrm{M}_{\odot}}$ for the upper bound. Stars with higher masses than this are already removed from the cluster mass accounting by stellar evolution in §2.1. We then evolve the mass function by removing stars until the sum of the remaining stellar mass is equal to the cluster mass at $z=0.3$. Since models of star cluster evolution (Baumgardt & Makino, 2003) show that low-mass stars tend to evaporate preferentially compared to high-mass stars, we remove stars probabilistically such that low-mass stars are more likely to be evaporated. Specifically, we treat the probability that any given star be removed as proportional to the inverse square root of its mass $m$:
@@
To further assess what effect the evaporation probability has on the stream profile, we selected a cluster and generated its stream twice, once with the adopted ejection probability $p\propto m^{-1/2}$ and once where every star is equally likely to be ejected independent of mass. After running the model, we binned the stars in the stream according to mass such that there were approximately 500 stars in each bin, and looked for variations between the two models as a function of mass. We found none significant. For example, the width of the stream, which we quantify with the standard deviation along the short axis of the stream, $\sigma_{\phi_{2}}$, shows little dependence on stellar mass: the best-fit slope of the linear regression was $0.16\pm 0.015$ for the stream with weighted ejection probability, and $0.010\pm 0.018$ for the stream with uniform ejection probability. Both of these slope are small, and we conclude that the evaporation probability has no significant effect on the structure of the stream and proceed to use it in our model to represent observations that older parts of streams have lower mass stars (Balbinot & Gieles, 2018).
@@
Figure 2: All sky map of the final collection of disrupted features for the TNG halo 523889 with $g<24.5$. Color indicates the metallicity of the original star cluster. Positions and apparent magnitudes of stars are calculated with the standard astropy convention where the Sun is placed at $(-8\rm\penalty 10000\ kpc,0\penalty 10000\ kpc,20.8\penalty 10000\ pc)$ in our galactocentric reference frame.
@@
where $\dot{M}(t)$ is the mass loss rate along the progenitor orbit and $m_{\rm av}(t)$ quantifies the average mass of ejected stars at time t. In terms of the ejection probability for a given star $p(m,t)$ (time-dependent because the normalization factor decreases as the number of stars in the cluster decreases) and the mass function of the cluster $\phi(m,t)$, $m_{\rm av}$ is equal to
@@
Figure 1 shows how the mass function of an example cluster in our model changes over time given the ejection probability of the individual stars. We take a cluster with initial mass $2\times 10^{4}\,{\mathrm{M}_{\odot}}$, a Kroupa IMF with lower and upper bounds of $0.1-1.1\,{\mathrm{M}_{\odot}}$, similar apocenter and pericenter parameters as Pal 5 (Odenkirchen et al., 2001), and integrate its orbit in a static version of one of our TNG potentials at the last snapshot for $13$ Gyr. We remove stars from the cluster given its mass loss rate from equation 2. At each age plotted, we update the upper mass bound based on the current location of the main sequence turn off, which varies slightly with cluster metallicity. For this figure we adopt a metallicity similar to Pal 5 of $\mathrm{[Fe/H]}=-1.56$ (Koch & Côté, 2017), which corresponds to a maximum mass of $0.8\,{\mathrm{M}_{\odot}}$ at present with the MIST isochrones. For the approximate metallicity range relevant to GCs, between 2.5-2.5 and 0.5-0.5 dex, the maximum mass at 13 Gyr ranges from $0.8-0.88\,{\mathrm{M}_{\odot}}$. After several Gyr, the slope of the low-mass end becomes less steep as low-mass stars are preferentially stripped from the cluster. This illustrates that streams realized with individual stars cannot be modeled by simply applying a zero-age IMF, because the cluster evolution significantly changes the mass distribution of the stream, especially for streams which are nearly or completely disrupted.
We assign stream particles numerically according to Equations 6 and 7. At each timestep, we take the mass-loss rate of the cluster and calculate how much mass needs to be removed. Then we remove 2 stars at a time (one for the leading tidal tail and one for the trailing, without replacement) from the bound stellar pool using the same method as described above, until the total mass released at a given timestep equals the appropriate amount according to the cluster mass-loss rate. This is repeated until the present day, giving us $N_{\rm stars}(t)$.
@@ 2.3.3 Orbit Integration
Given the number of stars released at each timestep, we integrate the stream orbits from when the stars are released (up to $\sim 3.5$ Gyr ago) until the present day. Using the Chen et al. (2025b) particle-spray method, stars are released near the Lagrange points, with their exact phase-space coordinates sampled from a multivariate gaussian distribution. Each cluster orbit is integrated in the combined potential of the galaxy and the cluster’s self-gravity, which is modeled as a Plummer sphere (Plummer, 1911). We continuously decrease the mass of the Plummer sphere according to the evolving mass of the cluster. Stars are released until the present day or until the cluster is fully disrupted, whichever occurs first. Once every stellar orbit has been completed we are left with the final stream resolved with individual stars at the present day. Repeating this process for each GC gives us the final stream catalog for each halo.
Figure 3: Four example streams with stars color-coded by their ejection lookback time. Each stream’s present day galactocentric radius is shown down and to the left of its progenitor location, as indicated by region where the ejection time approaches 0 Gyr.
Figure 4: Examples of streams from the catalog at a range of present day galactocentric radii with a photometric cut $g<24.5$. The top stream subtends the longest extent on the sky at nearly $250^{\circ}$ and its close proximity to the Galactic center has heated it and caused the stream to extend in $\phi_{2}$. The middle stream progenitor fully disrupted approximately $470\rm\penalty 10000\ Myr$ ago, causing a gap along the stream profile. The final stream at $47$ kpc from the Galactic center is dynamically the coldest, making it the shortest and thinnest of the three shown.
@@ 3 Results
In this section we calculate overall statistics for our catalog of disrupted GCs and make photometric predictions for present and upcoming missions. For plots we adopt the astropy convention and assume the Solar System to have galactocentric coordinates of $(-8\penalty 10000\ \mathrm{kpc},0,20.8\penalty 10000\ \mathrm{pc})$ in a right-handed system. Unless otherwise stated, we display only stars with an apparent magnitude $g<24.5$, the limit of a single visit for LSST (LSST Science Collaboration et al., 2009). All our photometry calculations are done with MIST Isochrones (Choi et al., 2016; Dotter, 2016). We also note that for our photometric predictions, we do not include extinction effects, and we perform our analysis across the entire sky.
@@ 3.1 Catalog Overview and Statistics
In total, our four halos (523889, 519311, m12i, m12w) contain 287, 401, 391, and 392 streams respectively. Figure 2 shows every stream for one of the TNG halos (523889) plotted on the sky with our fiducial magnitude cut. Each remnant is colored by its metallicity, which is the same value for each star of a given cluster. There is rich variation in the structure of tidal features across the entire population. Toward the Galactic center, there is a spherical grouping of stars from multiple clusters similar in morphology to the Galactic bulge (Wegg & Gerhard, 2013). These are predominantly clusters that have very small pericenters, where the tidal field is too strong to allow long, thin streams. At higher Galactic latitudes, there is an approximately uniform distribution of stars which have deviated significantly from their progenitor, identifiable as members of the stellar halo. Finally, there are many long and thin features which are clearly identifiable as stellar streams.
Figure 5: Variation of each stream in $(\phi_{1},\phi_{2})$ coordinates as a function of the galactocentric radius of the progenitor orbit for stars with $g<24.5$. The top row shows the dispersion for each stream in both $\phi_{1}$ and $\phi_{2}$, where as the bottom row shows the dispersion in the proper motions of each coordinate. Filled circles show thin streams, while open circles show wide streams. Streams closer to the Galactic center are dynamically hotter than streams further away and therefore exhibit a larger variation. Lines show best-fit power law relations at $r>10$ kpc for the three magnitude cuts: LSST (solid line), Roman (dot line), and Gaia (dot-dashed line).
@@
The average combined mass of all streams in each halo is $2\times 10^{7}\,{\mathrm{M}_{\odot}}$, ranging from $1.6\times 10^{7}\,{\mathrm{M}_{\odot}}$ (m12w) to $2.5\times 10^{7}\,{\mathrm{M}_{\odot}}$ (519311). The average total mass of stars in features within $3$ kpc of the Galactic center, roughly corresponding to the radius of the Galactic bulge, is $1.3\times 10^{7}\,{\mathrm{M}_{\odot}}$, ranging from $0.93\times 10^{7}\,{\mathrm{M}_{\odot}}$ (m12w) to $1.8\times 10^{7}\,{\mathrm{M}_{\odot}}$ (519311). Figure 3 shows four example thin streams from halo 523889, which span a range of present-day galactocentric radii 5205-20 kpc. Stars are colored by how long ago they were released from the cluster. A different set of streams spanning 5505-50 kpc is shown in Figure 4. For each stream in this Figure, the progenitor location is placed at $\phi_{1}=0^{\circ}$. The innermost stream, at a present day galactocentric radius of $5.4$ kpc, is the widest and has noticeable width variations along its track. The progenitor GC of the stream in the middle panel is fully disrupted at the present day, causing a gap to have formed in its center. There is also a slight asymmetry in the tidal tails caused by the projection into the $\phi_{1}-\phi_{2}$ frame. The outermost stream, at $47$ kpc, is the shortest, thinnest, and least massive, being subjected to the weakest tidal field out of the three. Collectively, these Figures illustrate the wide range of features and morphologies present in the overall MW stream population.
@@
Figure 5 shows how standard deviation in $\phi_{1}$, $\phi_{2}$, $\mu_{\phi_{1}}$ [^2] Here and in the rest of the paper, we use $\mu_{\phi_{1}}$ as shorthand for $\mu_{\phi_{1}}\cos{\phi_{2}}$., and $\mu_{\phi_{2}}$ vary as a function of distance to the Galactic center, as would be observed from Earth. Here, $\sigma_{\phi_{1}}$ characterizes the length of the stream, and $\sigma_{\phi_{2}}$ characterizes its width. These statistics are calculated for the stars in each stream that are below our fiducial magnitude cut, $g<24.5$.
@@
We distinguish here between what we consider wide streams (open circles) and thin streams (filled circles). These two groups are classified in the following way: first, we rotate each stream into galactocentric $(\phi_{1},\phi_{2})$ coordinates. This is done to emphasize the heating caused by an orbit being close to the Galactic center, and to de-emphasize observation effects, such as a stream with a blob-like morphology which appears thin due to its distance from the Earth. Any stream with $\sigma_{\phi_{2}}$ exceeding $10^{\circ}$ is considered a wide stream, which captures features that are too thick to be traditionally classified as a thin stream. Then we make a cumulative distribution of the ratio of a stream’s width to its length, $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}$, in Figure 6. Based on the flattening of the CDF gradient, we classify thin streams as having $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}<0.4$, corresponding to axis ratio greater than 2.5:1. We also repeated this procedure with an initial cut of $\sigma_{\phi_{s}}$ exceeding $5^{\circ}$ to check if our choice of initial cut significantly affects the CDF. As can be seen from Figure 6, there is a minimal difference between an initial cut of 5 or 10 degrees; for our classification we use 10 degrees.
@@
Beyond the inner galaxy, these relations all appear to follow a power law, but with significant intrinsic scatter. For each quantity $X$, we use maximum likelihood estimation to calculate the best-fitting power law relation of the form $X\propto r^{\alpha}$ for all points beyond $r>10$ kpc, because at smaller radii the relation visually appears to break down. This is influenced by the solar radius being at $\approx 8$ kpc, but we also find that the relation breaks around the same radius when the quantities in Figure 5 are calculated as seen from the Galactic center. To estimate the uncertainties for each slope, we resample our data $1000$ times and use the bootstrapping technique. For the standard deviation in $\phi_{1}$ and $\phi_{2}$ (in the top row of Figure 5) the best-fit slopes are $\alpha=-1.8\pm 0.3$ and $\alpha=-2.0\pm 0.5$. For the dispersion in proper motions along $\phi_{1}$ and $\phi_{2}$ (the bottom row of the figure), the slopes are $\alpha=-3.1\pm 0.6$ and $\alpha=-3.2\pm 0.5$.
@@
We also recompute what the slopes would be using magnitude cuts for Roman (dashed line) and Gaia (dot-dashed line). Details of how the cuts are made are in the next section (3.3). Overall the fits are mostly consistent between the three instruments. The width fit with the Gaia magnitude cut shows the greatest deviation from the other two, producing a slightly steeper ($\alpha=-2.2$) best-fit slope. This can be explained from the fact that the Gaia samples are most incomplete at large galactocentric radii. This suggests that width estimates for GC streams with members stars confirmed with Gaia may be underestimated.
@@
Figure 7: Cumulative distribution of the number of observable stars in each stream according to photometric completeness limits for Gaia ($G<19$), LSST ($g<24.5$), and Roman ($z<27.5$). Based on their reported magnitude limits, LSST and Roman will both drastically expand our ability to identify MW stellar streams.

### v2

@@ 2.3 Stellar Stream Generation
In this section, we describe the complete process for stream generation, given the Galactic potential and our initial GC conditions from the model of Chen and Gnedin (2024a). The distinct inputs to our stream model are as follows:
@@ 2.3.1 Mass Loss Calculation
The overall mass-loss calculation is in two parts. From cluster birth until $z=0.3$, we use the mass loss given by the GC formation model. Then beginning at $z=0.3$, we numerically solve Equation 2 given the tidal tensor along the cluster orbit computed with agama to calculate the amount of mass lost by the cluster from $z=0.3$ until the present day. While agama uses an adaptive timestep for orbit integration, we interpolate the solution such that the mass history and tidal tensor components are provided at 1000 uniformly spaced intervals over the integration time, giving us both quantities with a time resolution of $\sim 3.5$ Myr. Figure 12 in Appendix D shows the effective tidal frequency as a function of time at the provided resolution for an example cluster in one of the TNG halos. There are only four full TNG snapshots with potential information in our chosen time of integration, which miss significant variation of the tidal force. Consequently, the agama framework much more accurate for detailed calculations of the mass loss.
@@ 2.3.2 Stream Realization
Once calculated, the mass-loss histories are passed onto the next step, where we initialize a stellar population for each cluster and compute the number and masses of stream stars to be released at each timestep. First, individual cluster stars are sampled according to a Kroupa (2001) initial mass function (IMF) with the initial cluster mass given by the GC model. For the lower and upper bounds of the IMF, we use MIST isochrones (Choi et al., 2016; Dotter, 2016) though the isochrones interface (Morton, 2015) to calculate the theoretical lowest and highest mass stars that would still be luminous today given the cluster’s age and metallicity. These values vary from cluster to cluster but are generally around $0.1\,{\mathrm{M}_{\odot}}$ for the lower bound and between $0.8-0.9\,{\mathrm{M}_{\odot}}$ for the upper bound. Stars with higher masses than this are already removed from the cluster mass accounting by stellar evolution in §2.1. We then evolve the mass function by removing stars until the sum of the remaining stellar mass is equal to the cluster mass at $z=0.3$. Since models of star cluster evolution (Baumgardt and Makino, 2003) show that low-mass stars tend to evaporate preferentially compared to high-mass stars, we remove stars probabilistically such that low-mass stars are more likely to be evaporated. Specifically, we treat the probability that any given star be removed as proportional to the inverse square root of its mass $m$:
@@
To further assess what effect the evaporation probability has on the stream profile, we selected a cluster and generated its stream twice, once with the adopted ejection probability $p\propto m^{-1/2}$ and once where every star is equally likely to be ejected independent of mass. After running the model, we binned the stars in the stream according to mass such that there were approximately 500 stars in each bin, and looked for variations between the two models as a function of mass. We found none significant. For example, the width of the stream, which we quantify with the standard deviation along the short axis of the stream, $\sigma_{\phi_{2}}$, shows little dependence on stellar mass: the best-fit slope of the linear regression was $0.16\pm 0.015$ for the stream with weighted ejection probability, and $0.010\pm 0.018$ for the stream with uniform ejection probability. Both of these slopes are small, and we conclude that the evaporation probability has no significant effect on the structure of the stream. We proceed to use the probabilistic ejection setup in our model to represent observations that older parts of streams have lower mass stars (Balbinot and Gieles, 2018).
@@
Figure 2: All-sky map of the final collection of disrupted features for the TNG halo 523889 with $g<25$. Color indicates the metallicity of the original star cluster. Positions and apparent magnitudes of stars are calculated with the standard astropy convention where the Sun is placed at $(-8\rm\penalty 10000\ kpc,0\penalty 10000\ kpc,20.8\penalty 10000\ pc)$ in our galactocentric reference frame.
@@
where $\dot{M}(t)$ is the mass-loss rate along the progenitor orbit and $m_{\rm av}(t)$ quantifies the average mass of ejected stars at time t. In terms of the ejection probability for a given star $p(m,t)$ (time-dependent because the normalization factor decreases as the number of stars in the cluster decreases) and the mass function of the cluster $\phi(m,t)$, $m_{\rm av}$ is equal to
@@
Figure 1 shows how the mass function of an example cluster in our model changes over time given the ejection probability of the individual stars. We take a cluster with initial mass $2\times 10^{4}\,{\mathrm{M}_{\odot}}$, a Kroupa IMF with lower and upper bounds of $0.1-1.1\,{\mathrm{M}_{\odot}}$, similar apocenter and pericenter parameters as Pal 5 (Odenkirchen et al., 2001), and integrate its orbit in a static version of one of our TNG potentials (523889) at the last snapshot for $13$ Gyr. We remove stars from the cluster given its mass-loss rate from Equation 2. At each age plotted, we update the upper mass bound based on the current location of the main sequence turn off, which varies slightly with cluster metallicity. For this figure we adopt a metallicity similar to Pal 5 of $\mathrm{[Fe/H]}=-1.56$ (Koch and Côté, 2017), which corresponds to a maximum mass of $0.8\,{\mathrm{M}_{\odot}}$ at present with the MIST isochrones. For the approximate metallicity range relevant to GCs, roughly between 2.5-2.5 and 0.5-0.5 dex, the maximum mass at 13 Gyr ranges from $0.8-0.88\,{\mathrm{M}_{\odot}}$. After several Gyr, the slope of the low-mass end becomes less steep as low-mass stars are preferentially stripped from the cluster. This illustrates that streams realized with individual stars cannot be modeled by simply applying a zero-age IMF, because the cluster evolution significantly changes the mass distribution of the stream, especially for streams which are nearly or completely disrupted.
We assign stream particles numerically according to Equations 6 and 7. At each timestep, we take the mass-loss rate of the cluster and calculate how much mass needs to be removed. Then, we remove 2 stars at a time (one for the leading tidal tail and one for the trailing, without replacement) from the bound stellar pool using the same method as described above, until the total mass released at a given timestep equals the appropriate amount according to the cluster mass-loss rate. This is repeated until the present day, giving us $N_{\rm stars}(t)$.
@@ 2.3.3 Orbit Integration
Given the number of stars released at each timestep, we integrate the stream orbits from when the stars are released (up to $\sim 3.5$ Gyr ago) until the present day. Using the Chen et al. (2025b) particle-spray method, stars are released near the Lagrange points, with their exact phase-space coordinates sampled from a multivariate gaussian distribution. This phase space distribution utilized by the Chen et al. (2025b) algorithm was tested in static, axisymmetric Milky Way potentials, which are inherently less complex than the potentials used in this work, which are both time-evolving and non-axisymmetric. In Appendix E, we do a comparison with one stream from our catalog generated with full N-body, and verify that the Chen et al. (2025b) shows excellent agreement with the N-body realization even in our time-evolving, non-axisymmetric regime.
Each orbit is integrated using our particle-spray model in the combined potential of the galaxy and the cluster’s self-gravity, which is modeled as a Plummer sphere (Plummer, 1911). We continuously decrease the mass of the Plummer sphere according to the evolving mass of the cluster. We set the Plummer scale radius of each cluster to $4\penalty 10000\ \mathrm{pc}$. While this value should decrease as the cluster loses mass, its exact value will not effect the orbit of escaped stars if the cluster’s tidal radius is much greater than the scale radius. For our clusters that do not fully disrupt, we find this to be true in almost all cases; for fully disrupted clusters, we find this to be true down until the cluster mass is in the range $10^{2}-10^{3}\,{\mathrm{M}_{\odot}}$. As such, we keep the scale radius fixed for simplicity. Stars are released until the present day or until the cluster is fully disrupted, whichever occurs first. Once every stellar orbit has been completed we are left with the final stream resolved with individual stars at the present day. Repeating this process for each GC gives us the final stream catalog for each halo.
Figure 3: Four example streams with stars color-coded by their ejection lookback time. Each stream’s present day galactocentric radius is shown down and to the left of its progenitor location, as indicated by region where the ejection time approaches either 0 Gyr or the time of disruption.
Figure 4: Examples of streams from the catalog at a range of present day galactocentric radii with a photometric cut $g<25$. The top stream subtends the longest extent on the sky at nearly $250^{\circ}$ and its close proximity to the Galactic center has heated it and caused the stream to extend in $\phi_{2}$. The middle stream progenitor fully disrupted approximately $470\rm\penalty 10000\ Myr$ ago, causing a gap along the stream profile. The final stream at $47$ kpc from the Galactic center is dynamically the coldest, making it the shortest and thinnest of the three shown.
@@ 3 Results
In this section we calculate overall statistics for our catalog of disrupted GCs and make photometric predictions for present and upcoming missions. For plots we adopt the astropy convention and assume the Solar System to have galactocentric coordinates of $(-8\penalty 10000\ \mathrm{kpc},0,20.8\penalty 10000\ \mathrm{pc})$ in a right-handed system. Unless otherwise stated, we display only stars with an apparent magnitude $g<25$, the limit of a single visit for LSST (LSST Science Collaboration et al., 2009). All our photometry calculations are done with MIST Isochrones (Choi et al., 2016; Dotter, 2016), and stream properties are calculated only using stars below our photometric magnitude cuts. We also note that for our photometric predictions, we do not include extinction effects.
@@ 3.1 Catalog Overview and Statistics
In total, our four halos (523889, 519311, m12i, m12w) contain 287, 401, 391, and 392 streams respectively. Table 1 shows for each halo the total number of streams, the total mass of stars in streams, as well as the total number of stars in streams. Figure 2 shows every stream for one of the TNG halos (523889) plotted on the sky with our fiducial magnitude cut. Each remnant is colored by its metallicity, which is the same value for each star of a given cluster. There is rich variation in the structure of tidal features across the entire population. Toward the Galactic center, there is a spherical grouping of stars from multiple clusters similar in morphology to the Galactic bulge (Wegg and Gerhard, 2013). These are predominantly clusters that have very small pericenters, where the tidal field is too strong to allow long, thin streams. At higher Galactic latitudes, there is an approximately uniform distribution of stars which have deviated significantly from their progenitor, identifiable as members of the stellar halo. Finally, there are many long and thin features which are clearly identifiable as stellar streams.
Figure 5: Variation of each stream in $(\phi_{1},\phi_{2})$ coordinates across all four halos as a function of the galactocentric radius of the progenitor orbit for stars with $g<25$. The top row shows the dispersion for each stream in both $\phi_{1}$ and $\phi_{2}$, where as the bottom row shows the dispersion in the proper motions of each coordinate. Filled circles show thin streams, while open circles show wide streams. Streams closer to the Galactic center are dynamically hotter than streams further away and therefore exhibit a larger variation. Lines show best-fit power law relations at $r>10$ kpc for the three magnitude cuts: LSST (solid line), Roman (dot line), and Gaia (dot-dashed line).
@@
The average combined mass of all streams in each halo is $2\times 10^{7}\,{\mathrm{M}_{\odot}}$, ranging from $1.6\times 10^{7}\,{\mathrm{M}_{\odot}}$ (m12w) to $2.5\times 10^{7}\,{\mathrm{M}_{\odot}}$ (519311). The average total mass of stars in features within $3$ kpc of the Galactic center, roughly corresponding to the radius of the Galactic bulge, is $1.3\times 10^{7}\,{\mathrm{M}_{\odot}}$, ranging from $0.93\times 10^{7}\,{\mathrm{M}_{\odot}}$ (m12w) to $1.8\times 10^{7}\,{\mathrm{M}_{\odot}}$ (519311). Figure 3 shows four example thin streams from halo 523889, which span a range of present-day progenitor galactocentric radii 5205-20 kpc. Stars are colored by how long ago they were released from the cluster. A different set of streams spanning 5505-50 kpc is shown in Figure 4. For each stream in this Figure, the progenitor location is placed at $\phi_{1}=0^{\circ}$. The innermost stream, at a present day galactocentric radius of $5.4$ kpc, is the widest and has noticeable width variations along its track. The progenitor GC of the stream in the middle panel is fully disrupted at the present day, causing a gap to have formed in its center. There is also a slight asymmetry in the tidal tails caused by the projection into the $\phi_{1}-\phi_{2}$ frame. The outermost stream, at $47$ kpc, is the shortest, thinnest, and least massive, being subjected to the weakest tidal field out of the three. Collectively, these Figures illustrate the wide range of features and morphologies present in the overall MW stream population.
Table 1: Overall statistics for each halo including the number of streams, total number of stars in streams, and total mass of streams for each halo.
Halo ID Nstreams N[$_{\rm stars}\penalty 10000\ [10^{7}$] M[$_{\rm streams}\penalty 10000\ [10^{7}\,{\mathrm{M}_{\odot}}$]
523889 $287$ $6.1$ $2.0$
519311 $401$ $7.8$ $2.5$
m12i $391$ $5.4$ $1.7$
m12w $392$ $5.1$ $1.6$
@@
Figure 5 shows how standard deviation in $\phi_{1}$, $\phi_{2}$, $\mu_{\phi_{1}}$ [^2] Here and in the rest of the paper, we use $\mu_{\phi_{1}}$ as shorthand for $\mu_{\phi_{1}}\cos{\phi_{2}}$., and $\mu_{\phi_{2}}$ vary as a function of distance to the Galactic center, as would be observed from Earth. Here, $\sigma_{\phi_{1}}$ characterizes the length of the stream, and $\sigma_{\phi_{2}}$ characterizes its width. For the proper motion dispersions and stream length, we calculate each quantity by rotating into the stream frame and computing the standard deviation of $\mu_{\phi_{1}},\mu_{\phi_{2}},$ and $\phi_{1}$. For the width of the stream, we first use agama to fit a spline to the stream track in $(\phi_{1},\phi_{2})$ coordinates and calculate the standard deviation from the fitted track. These statistics are calculated only using the stars in each stream that are below our fiducial magnitude cut, $g<25$.
@@
We distinguish here between what we consider wide streams (open circles) and thin streams (filled circles). While wide streams often in the literature refer to dwarf galaxy streams, here we take wide streams to be GC streams which have undergone significant heating due to their orbit. These two groups are classified in the following way: first, we rotate each stream into galactocentric $(\phi_{1},\phi_{2})$ coordinates. This is done to emphasize the heating caused by an orbit being close to the Galactic center, and to de-emphasize observation effects, such as a stream with a blob-like morphology which appears thin due to its distance from the Earth. Any stream with $\sigma_{\phi_{2}}$ exceeding $10^{\circ}$ is considered a wide stream, which captures features that are too thick to be traditionally classified as a thin stream. Then we make a cumulative distribution of the ratio of a stream’s width to its length, $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}$, in Figure 6. Based on the flattening of the CDF gradient, we classify thin streams as having $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}<0.4$, corresponding to axis ratio greater than 2.5:1. We also repeated this procedure with an initial cut of $\sigma_{\phi_{s}}$ exceeding $5^{\circ}$ to check if our choice of initial cut significantly affects the CDF. As can be seen from Figure 6, there is a minimal difference between an initial cut of 5 or 10 degrees; for our classification we use 10 degrees.
Table 2: Power Law fits for the streams as a function of galactocentric distance
Instrument $\sigma_{\phi_{1}}$ $\sigma_{\phi_{2}}$ $\sigma_{\mu_{\phi_{1}}}$ $\sigma_{\mu_{\phi_{2}}}$
LSST $1.8\pm{0.3}$ $1.9\pm{0.6}$ $3.0\pm{0.5}$ $3.2\pm{0.5}$
Roman $1.8\pm{0.3}$ $1.7\pm{0.6}$ $2.9\pm{0.5}$ $3.0\pm{0.5}$
Gaia $1.3\pm{0.4}$ $2.8\pm{0.9}$ $1.9\pm{0.7}$ $2.0\pm{0.7}$
@@
Beyond the inner galaxy, these relations all appear to follow a power law, but with significant intrinsic scatter. For each quantity $X$, we use maximum likelihood estimation to calculate the best-fitting power law relation of the form $X\propto r^{\alpha}$ for all points beyond $r>10$ kpc, because at smaller radii the relation visually appears to break down. This is influenced by the solar radius being at $\approx 8$ kpc, but we also find that the relation breaks around the same radius when the quantities in Figure 5 are calculated as seen from the Galactic center. To estimate the uncertainties for each slope, we resample our data $1000$ times and use the bootstrapping technique. For the standard deviation in $\phi_{1}$ and $\phi_{2}$ (in the top row of Figure 5) the best-fit slopes are $\alpha=-1.8\pm 0.3$ and $\alpha=-1.9\pm 0.6$. For the dispersion in proper motions along $\phi_{1}$ and $\phi_{2}$ (the bottom row of the figure), the slopes are $\alpha=-3.0\pm 0.5$ and $\alpha=-3.2\pm 0.5$.
@@
We also recompute what the slopes would be using magnitude cuts for Roman (dashed line) and Gaia (dot-dashed line). Details of how the cuts are made are in the next section (3.3). We show the full list of best-fit slopes for each instrument in Table 2. Overall, both LSST and Roman are consistent with each other, with the prediction from Gaia varying significantly in comparison. This can be explained from the fact that the Gaia samples are mostly incomplete at large galactocentric radii, with only a few observable stars in streams beyond $30$ kpc. As a result, the width relation from Gaia is underestimated at large galactocentric radii, whereas the other three quantities are overestimated. Most significant of these results is that width estimates for GC streams with member stars confirmed with Gaia may be underestimated. In addition to the scatter already present in the width distribution, it is likely that variation in GC concentration (which is not captured in this work) would add additional scatter.
@@
Figure 7: Cumulative distribution of the number of streams with $N_{\rm stars}$ or more observable stars according to photometric completeness limits for Gaia ($G<19$), LSST ($g<25$), and Roman ($z<27.5$). Based on their reported magnitude limits, LSST and Roman will both drastically expand our ability to identify MW stellar streams. Dashed line shows the total number of streams.